Phantomdex:

The probability of event A occurring is 0.27, the probability of event B occurring is 0.19, and the probability of events A and B occurring is 0.11. Determine P(A|B). 0.051 0.350 0.407 0.579

Phantomdex:

@toga

toga:

To determine P(A|B), we can use the formula: P(A|B) = P(A and B) / P(B) We are given that P(A) = 0.27, P(B) = 0.19, and P(A and B) = 0.11. Using the formula, we get: P(A|B) = 0.11 / 0.19 Simplifying, we get: P(A|B) = 0.579 Therefore, the probability of event A occurring given that event B has occurred is 0.579.